A car moving with a velocity of 54 km/h accelerates uniformly at the rate of 2 ms-2. Calculate the distance travelled from the place where acceleration began to that where the velocity reaches 73km/h. And the time taken to cover this distance.


1
Expert's answer
2021-03-02T18:03:10-0500

"s=\frac{v^2-v_0^2}{2a}=\frac{20.28^2-15^2}{2\cdot 2}=46.57\ (m)" . Answer


"v=v_0+at\to t=(v-v_0)/a=(20.28-15)/2=2.64\ (s)" . Answer